# A Comprehensive Evaluation on Event Reasoning of Large Language Models

### v1

@@ Abstract
Event reasoning is a fundamental ability that underlies many applications. It requires event schema knowledge to perform global reasoning and needs to deal with the diversity of the inter-event relations and the reasoning paradigms. How well LLMs accomplish event reasoning on various relations and reasoning paradigms remains unknown. To mitigate this disparity, we comprehensively evaluate the abilities of event reasoning of LLMs. We introduce a novel benchmark for EValuation of EVent reasoning. consists of two levels of evaluation of schema and instance and is comprehensive in relations and reasoning paradigms. We conduct extensive experiments on . We find that LLMs have abilities to accomplish event reasoning but their performances are far from satisfactory. We also notice the imbalance of event reasoning abilities in LLMs. Besides, LLMs have event schema knowledge, however, they’re not aligned with humans on how to utilize the knowledge. Based on these findings, we introduce two methods to guide the LLMs to utilize the event schema knowledge. Both methods achieve improvements. Code and Dataset are available on https://github.com/TZWwww/EV2.
@@ A Comprehensive Evaluation on Event Reasoning of Large
Zhengwei Tao12 Zhi Jin12††thanks: *Corresponding authors. Yifan Zhang12 Xiancai Chen12 Xiaoying Bai3∗ Yue Fang12 Haiyan Zhao12 Jia Li12 Chongyang Tao4 1Key Laboratory of High Confidence Software Technologies (PKU), MOE, China 2School of Computer Science, Peking University 3Advanced Institute of Big Data 4Beihang University {tttzw, xiancaich, yifanzhang, y.fang}@stu.pku.edu.cn, baixy@aibd.ac.cn {zhijin,zhhy.sei, lijiaa}@pku.edu.cn chongyang@buaa.edu.cn
@@ 1 Introduction
Events are instances or occurrences that form the basic semantic building units encompassing the meanings of Activities, Accomplishments, Achievements, and States Vendler (1957). Event Reasoning is the ability to process and analyze events and their complex interconnections. Compared with other abilities, event reasoning is unique in some aspects. Firstly, it requires knowledge in the form of event schemas, capturing the progress of event evolution in scenarios, then performing global reasoning Li et al. (2021a); Mao et al. (2021). As shown in Figure 1, each event instance is associated with an event type. All event types and their relations form the event schema knowledge which reflects the logic and mechanism of event evolution. Knowing “Memory” would often happen after “Learn” can help answer the reasoning question. Second, the inter-event relations and reasoning paradigms are various. Event reasoning incorporates reasoning events according to a certain relation Du et al. (2022); Sap et al. (2019b) and reasoning inter-event relations Ning et al. (2018); Caselli and Vossen (2017). The queried relations are diversified such as causality Roemmele et al. (2011), temporality Zhou et al. (2019), and hierachy Glavaš et al. (2014). There are various paradigms such as reasoning the event or the inter-relation.
As a fundamental competency within LLMs, event reasoning supports a multitude of Natural Language Processing (NLP) tasks, including recommendation engines Yang et al. (2020), interactive question-answer systems Souza Costa et al. (2020), and AI Agents Liu et al. (2023). Therefore, the enhancement of event reasoning abilities is essential for the advancement of LLMs.
LLMs like LLAMA Touvron et al. (2023) series and GPT series Brown et al. (2020) have demonstrated exceptional accomplishments in various natural language reasoning Bang et al. (2023); Xu et al. (2023b). Existing research has evaluated a broad spectrum of reasoning abilities of LLMs such as commonsence Bian et al. (2023), sentence relations Chan et al. (2023), and math Arora et al. (2023). However, studies on the comprehensive evaluation of event reasoning of LLMs are scarce. Current works only focus on instance-level events, resulting in unclearness of how LLMs understand and utilize the event schema knowledge Chan et al. (2023). Besides, they ignore the diversity of relations and paradigms Yuan et al. (2023). These disparities hinge on the development of such crucial abilities of LLMs.
@@
We conduct extensive experiments on to answer these questions. The results provide insights into event reasoning that: 1) LLMs have the abilities of event reasoning, but are far from satisfactory and are imbalanced in different relations and reasoning paradigms. 2) LLMs have event schema knowledge. They can answer the schema-level questions with similar accuracy to the instance-level questions. However, the development of schema-level abilities falls behind those of instance-level. 3) LLMs are not aligned with humans in the aspect of leveraging event schema knowledge. 4) Based on the findings, we design two mentoring methods to guide the LLMs to utilize event schema knowledge. One is to directly add event schema knowledge to the prompt. The second is guiding in a chain-of-thought format. With the designed guidances for utilizing event schema knowledge, LLMs can perform better event reasoning. Especially with direct guidance, LLMs get significant improvements.
@@
We conclude several insights. Based on our findings, we design mentoring methods to guide LLMs to utilize event schema knowledge which achieves improvements in event reasoning.
@@ 2 Problem Formulation
Event reasoning is to anticipate the occurrences of certain relations or deduce interrelated correlations Tao et al. (2023a). The relations encompass causality Du et al. (2022), temporality Zhou et al. (2019), and hierarchy Glavaš et al. (2014).
Event reasoning requires comprehension of event schema knowledge. An event schema of a scenario is a schema-level graph 111Superscript represents schema level., where is the set of event types and is the set of relations between events. Each edge in is a relation triplet standing for that there is the relation between and . With instantiation, we have the instance-level event graph 222Superscript represents instance level.. An instance event has an event type but with detailed event arguments and context Mitchell (2005). The nodes and edges of these two graphs are corresponding, namely, each triplet in has a corresponding triplet in with the same inter-relation. In both levels, we consider totally six relation types, namely {Causes, IsResult, Before, After, IsSubevent, HasSubevent}.
@@ 3 Benchmark Construction
To create the benchmark, we curate a comprehensive dataset through a three-stage process. Initially, the schema graph is established. Then, GPT4 is employed to generate the instance graph . Lastly, human annotators are tasked with creating questions from and .
3.1 Schema Graph Construction
We leverage EECKG Wang et al. (2022b) to ensure a diverse range of event types in our schema. EECKG combines rule-based reasoning with crowdsourced insights, built on ConceptNet’s structure. Nodes in EECKG represent verb phrases as events, and edges denote inter-event relations, focusing on Causes333The direction is that the head event causes the tail event. Other relations are the same., Before, and HasSubevent.
@@ 3.2 Instance Graph Construction
We inherit the relations of and obtain . We naturally obtain the instances of and .
3.3 Question Construction
The last step is to construct questions of CEC and CRR in both schema and instance levels. For CEC, regarding schema and instance head events as the query and the tail as an answer, we ask GPT4 to generate 15 possible candidate instance events with their event types.
We then recruit 8 well-educated human annotators. Their missions are:
@@ 3.3 Question Construction
Revise or discard if not valid. Ensure the events are abstract, the relations are correct, there’s no scenario shifting in .
@@
Revise or discard if not valid. Ensure the events are concrete, the relations are correct, and the whole scenario of is coherent and has no shifting.
@@
Choose three proper negative candidate events with their event types. Ensure answering the question should rely on the context events.
We use the schema part of annotation as the schema-level questions and the instance part as instance-level questions. Then we complete CEC.
For CRR, we regard and as queried events and use the relation between them as the answer to form the schema-level question. For instance part, we adopt a similar way.
Our CEC task is a 4-way multiple-choice task. The CRR is a 3-way multiple-choice task. In CRR, the choices for temporal, causal, and hierarchy relations are [Before, After, Vague], [Causes, IsResult, None], and [IsSubevent, HasSubevent, None] respectively. We show examples of both tasks in Figure 1. We report the number of each task and the average nodes and edges of in Figure 2.
@@
| S-CEC | I-CEC | S-CRR | S-CRR | AVG N | AVG E |
| 492 | 558 | 767 | 835 | 3.62 | 2.78 |
@@ 3.4 Quality Inspection
We recruit other human annotators to inspect the quality of . We sample 100 data for all tasks. We ask them to give two scores for each sample:
@@ 4.1 Evaluated LLMs
We evaluate 9 LLMs on event reasoning. For the open-source models, we evaluate their chat-version. For the closed-source models, we utilize their official APIs to conduct performance evaluations. Specifically, we employ the gpt-4-0125-preview version as the GPT4 and the gpt-3.5-turbo-1106 version as GPT3.5 in our experiments. For the open-source models, we include Qwen1.5-7B Bai et al. (2023), Mistral-7B Jiang et al. (2023), Baichuan-2-7B Yang et al. (2023), Llama2-7B Touvron et al. (2023), WizardLM-7B Xu et al. (2023a), Vicuna-7B Chiang et al. (2023), and Alpaca-7B Taori et al. (2023). Without loss of generosity, we use the model names to refer to the chat versions in the rest of our paper. For all evaluated LLMs, we use the same prompt. We show prompts in Figure 5-8 in the Appendix.
@@ 5.1 How proficient abilities of event reasoning do LLMs have?
We evaluate CEC and CRR at the instance level. We show the results of different relations in Figure 2. For CEC, GPT4 performs the best. Models like Qwen1.5-7B, Mistral-7B, and GPT3.5 are in the second tier. Qwen1.5-7B and Mistral-7B are both better than GPT3.5. Qwen1.5-7B can even excel GPT4 in the temporal and causal relations. The other models such as WizardLM-7B almost fail, obtaining lower than 40% accuracy. For CRR, GPT4 excels all other models as well. However, unlike CEC, there is no obvious difference in the performance of other models for CRR.
We show the average performance of instance-level CEC and CRR in columns I-CEC and I-CRR in Table 4. We only show models that have basic abilities, namely CEC accuracy above 50.00 or CRR above 40.00 in Table 4, while other models may lack analytical significance. Overall, existing LLMs such as GPT4, and Qwen1.5-7B have CECtain event reasoning abilities. However, even the strongest GPT4 can only achieve 63.80 and 61.20 accuracy in each task showing there’s much room for improvements of event reasoning.
@@
Comparing CEC to CRR, as relation-wise results shown in Figure 2 and average performances in columns I-CEC and I-CRR in Table 4, LLMs perform better for CEC than CRR. We compute the average scores of four listed models in Table 4. We find I-CEC is much higher than I-CRR, with 58.91 to 46.18. The results significantly suggest that CRR is harder than CEC. Existing pretraining and SFT datasets may be biased in paradigms.
We then analyze performances on different relations. As shown in Figure 2, LLMs perform best in causality relation. Then, temporal, and hierarchy relations are tied. That further indicates the imbalance training of different relations. Methods and datasets of balanced abilities on relations are needed. Transferring abilities of different relations could also be feasible Tao et al. (2023b).
This is a crucial finding. Chan et al. (2023) conduct causal event classification such as ECARE Du et al. (2022), and relation reasoning such as MATRES Ning et al. (2018). They directly compare these two groups of results and conclude the gaps are merely from differences in relations. However, they ignore the difference in reasoning paradigms. Leveraging , with disentangling relations and formulations, we investigate event reasoning with less bias.
LLMs excel in forward CEC compared with backward.
We calculate the average scores of forward relation (After, IsResult) and backward relation (Before, Causes). The results are shown in columns I-F, and I-B in Table 4. We find that the average of I-F is significantly better than I-B. It also suggests that the training dataset is unbalanced in relations. Less of the training data is used for backward relations, resulting in poorer performances on those. However, backward relations are important in abduction scenarios. Methods should be designed to enhance such abilities.
@@
We investigate the improvement trends of CEC and CRR. In Figure 3. When models have poor event reasoning abilities, their performances lie around the balanced line showing no significant differences in tasks. With the development, the CEC improves much faster than CRR such models as GPT3.5, Mistral-7B, and Qwen1.5-7B. This investigation appeals to the need for training in comprehensive event reasoning abilities.
| Model | S-F | S-B | I-F | I-B | S-CEC | I-CEC | S-CRR | I-CRR |
| GPT4 | 54.22 | 55.84 | 65.34 | 61.84 | 55.48 | 63.80 | 52.80 | 61.20 |
| GPT3.5 | 47.88 | 50.94 | 55.68 | 49.12 | 49.79 | 50.18 | 45.37 | 39.52 |
| Qwen | 42.96 | 52.45 | 67.05 | 65.37 | 48.98 | 63.98 | 43.00 | 40.00 |
| Mistral | 44.37 | 52.08 | 67.61 | 53.71 | 48.98 | 57.71 | 46.00 | 44.00 |
| AVG | 47.35 | 52.82 | 63.92 | 57.51 | 50.8 | 58.91 | 46.79 | 46.18 |

### v2

@@ Abstract
Event reasoning is a fundamental ability that underlies many applications. It requires event schema knowledge to perform global reasoning and needs to deal with the diversity of the inter-event relations and the reasoning paradigms. How well LLMs accomplish event reasoning on various relations and reasoning paradigms remains unknown. To mitigate this disparity, we comprehensively evaluate the abilities of event reasoning of LLMs. We introduce a novel benchmark for EValuation of EVent reasoning. consists of two levels of evaluation of schema and instance and is comprehensive in relations and reasoning paradigms. We conduct extensive experiments on . We find that LLMs have abilities to accomplish event reasoning but their performances are far from satisfactory. We also notice the imbalance of event reasoning abilities in LLMs. Besides, LLMs have event schema knowledge, however, they’re not aligned with humans on how to utilize the knowledge. Based on these findings, we guide the LLMs in utilizing the event schema knowledge as memory leading to improvements on event reasoning. Code and Dataset are available on https://github.com/TZWwww/EV2.
@@ A Comprehensive Evaluation on Event Reasoning of Large
Zhengwei Tao12 Zhi Jin1††thanks: *Corresponding author. Yifan Zhang1 Xiancai Chen1 Haiyan Zhao1 Jia Li1 Bing Liang2 Chongyang Tao3 Qun Liu4 Kam-Fai Wong2 1Key Lab of HCST (PKU), MOE; School of Computer Science, Peking University 2Department of Systems Engineering and Engineering Management, The Chinese University of Hong Kong 3Beihang University 4Huawei Noah’s Ark Lab {tttzw, xiancaich, yifanzhang}@stu.pku.edu.cn {zhijin, zhhy.sei, lijiaa}@pku.edu.cn, chongyang@buaa.edu.cn bin.liang@cuhk.edu.hk, qun.liu@huawei.com, kfwong@se.cuhk.edu.hk
@@ 1 Introduction
Events are instances or occurrences that form the basic semantic building units encompassing the meanings of Activities, Accomplishments, Achievements, and States Vendler (1957). Event Reasoning is the ability to process and analyze events and their complex interconnections. Compared with other abilities, event reasoning is unique in some aspects. Firstly, it requires knowledge in the form of event schemas, capturing the progress of event evolution in scenarios, then performing global reasoning Li et al. (2021a); Mao et al. (2021). As shown in Figure 1, each event instance is associated with an event type. All event types and their relations form the event schema knowledge which reflects the logic and mechanism of event evolution. Knowing the event occurrence chain as "Avoiding", "Enjoy life", "Celebrate", and "Hold Party" would result in the following event "Express Emotions". Second, the inter-event relations and reasoning paradigms are various. Event reasoning incorporates reasoning events according to a certain relation Du et al. (2022a); Sap et al. (2019b) and reasoning inter-event relations Ning et al. (2018); Caselli and Vossen (2017). The queried relations are diversified such as causality Roemmele et al. (2011), temporality Zhou et al. (2019), and hierachy Glavaš et al. (2014). There are various paradigms such as reasoning the event or the inter-relation. As a fundamental competency within LLMs, event reasoning supports a multitude of Natural Language Processing (NLP) tasks, including recommendation engines Yang et al. (2020), interactive question-answer systems Souza Costa et al. (2020), and AI Agents Liu et al. (2023). Therefore, the enhancement of event reasoning abilities is essential for the advancement of LLMs.
LLMs like LLAMA Touvron et al. (2023) series and GPT series Brown et al. (2020) have demonstrated exceptional accomplishments in various natural language reasoning Bang et al. (2023); Xu et al. (2023). Existing research has evaluated a broad spectrum of reasoning abilities of LLMs such as commonsence Bian et al. (2023), sentence relations Chan et al. (2023), and math Arora et al. (2023). However, studies on the comprehensive evaluation of event reasoning of LLMs are scarce. The incompleteness of current event reasoning evaluation is reflected in two aspects. First, current works only focus on instance-level events, resulting in unclearness of how LLMs understand and utilize the event schema knowledge Chan et al. (2023). Investigating the event knowledge of LLMs and how they employ them underlines applications such as event-based memory systems. Besides, they ignore the diversity of relations and paradigms Yuan et al. (2023). Such findings could be biased since they neglect discrepancies brought by different aspects. These disparities hinge on the development of such crucial abilities of LLMs.
@@
We conduct extensive experiments on to answer these questions. The results provide insights into event reasoning that: 1) LLMs have the abilities of event reasoning, but are far from satisfactory and are imbalanced in different relations and reasoning paradigms. 2) LLMs have event schema knowledge. They can answer the schema-level questions with similar accuracy to the instance-level questions. However, the development of schema-level abilities falls behind those of instance-level. 3) LLMs are not aligned with humans in the aspect of leveraging event schema knowledge. 4) Based on the findings, we investigate guiding the LLMs to utilize event schema knowledge. With the guidance, LLMs can perform better event reasoning which sheds light on modeling event knowledge as memory of LLMs to enhance event reasoning.
@@
We conclude several insights. Based on our findings, we guide LLMs to utilize event schema knowledge as memory achieving improvements in event reasoning.
@@ 2 Problem Formulation
Event reasoning is to anticipate the occurrences of certain relations or deduce interrelated correlations Tao et al. (2023a). The relations encompass causality Du et al. (2022a), temporality Zhou et al. (2019), and hierarchy Glavaš et al. (2014).
Event reasoning requires comprehension of event schema knowledge. An event schema of a scenario is a schema-level graph 111Superscript represents schema level., where is the set of event types and is the set of relations between events. Each edge in is a relation triplet standing for that there is the relation between and . With instantiation, we have the instance-level event graph 222Superscript represents instance level.. An instance event has an event type but with detailed event arguments and context Mitchell (2005). The nodes and edges of these two graphs are corresponding, namely, each triplet in has a corresponding triplet in with the same inter-relation. In both levels, we consider totally six relation types, namely {Causes, IsResult, Before, After, IsSubevent, HasSubevent}333In this work, we denote the direction of the relation to that the former event have relation to the latter event. For example, (, Causes, ) stands for causes .. Causes and IsResult are Causal relations, Before and After belong to Temporal type while IsSubevent and HasSubevent are Hierarchical type.
@@ 3 Benchmark Construction
Constructing the benchmark is challenging since events and their relations are semantically abstract concepts compared with entity concepts. The occurrence of events and their relations not only follow objective natural laws but are also influenced by social and humanistic factors. Therefore, annotating such data is extremely label-intensive leading to a lack of evaluation dataset of the task. Previous works mainly construct such datasets by extracting events and relations from some unlabeled corpus such as news reports Caselli and Vossen (2017); Ning et al. (2018); O’Gorman et al. (2016). However, such a method suffers from limited event relational patterns and domain-specific language expression. To mitigate such problems, in this work, we construct our comprehensive event reasoning evaluation from scratch.
In , we evaluate event reasoning in various relation and reasoning paradigms of both schema and instance levels. In our pilot trial, we directly annotate each question and answer by human annotators. We find this process extremely time-consuming since human annotators must imagine scenarios and guarantee their correctness. Besides, the diversity of the events and their relations are poor. Annotators only write the most common events. Therefore, we propose a four-stage construction process. Generally, we first synthesize both the schema and corresponding instance event prompting-graphs automatically as prompts for later annotations. Due to the large size of the synthesis graphs, before formal annotating, we then recruit annotators to remove incorrect graphs which are hard to modify. After that, the annotators curate schema and instance graphs based on the generated prompts. Finally, we adapt the graphs into questions and answers. We celebrate this process in the following sections.
3.1 Prompting-Graph Synthesizing
To cover more domain and granular scenarios, we synthesize event graphs as prompts for annotations instead of annotating from scratch. Specifically, we first establish the schema graph . Then we employ GPT4 to generate the instance graph .
Schema Graph
The schema graph represents event occurrence knowledge. Harvesting such knowledge is a research point Du et al. (2022b). We here leverage EECKG Wang et al. (2022b) to ensure a diverse range of event types in our schema. EECKG combines rule-based reasoning with crowdsourced insights, built on ConceptNet’s structure. Nodes in EECKG represent verb phrases as events, and edges denote inter-event relations, focusing on Causes, Before, and HasSubevent.
@@ 3.1 Prompting-Graph Synthesizing
We inherit the relations of and obtain . We naturally obtain the instances of and . Finally, we obtain 1,600 schema prompting graphs and 1,600 corresponding instance graphs.
3.2 Manual Filtering
After curating the prompting graphs of both schema and instance levels, the next is to annotate based on the prompting graphs. However, we find some of the prompting graphs are incorrect and hard to modify. Therefore, before formal annotation, we launch another manual filtering step to remove such graphs.
We then recruit 8 well-educated human annotators where they each process 200 data. Their missions are to investigate the and by the following steps:
@@ 3.2 Manual Filtering
Check whether can be modified, the events are abstract, the relations are correct, and tells an entire story of a scenario.
@@
Check whether can be modified., the events are concrete, the relations are correct, and tells an entire story of a scenario.
@@
Check whether and are consistent, there are obvious schema-instance relations between them.
If any of these conditions are not met, we discard this datum. After this filtering process, we remain 491 graph pairs.
@@
S-CEC | I-CEC | S-CRR | I-CRR | Graph Pairs |
492 | 491 | 730 | 735 | 491 |
3.3 Annotation
In this stage, we formally annotate based on filtered prompting-graphs. The missions of this stage are to 1) rewrite the events and relations in both schema and instance graphs to make them strictly valid. 2) identify a query event and an answer event in the graph for later question adaptation. 3) write candidates as negative choices considering the answer event where each candidate event consists of a schema and an instance event. For the second mission, regarding schema and instance head events as the query and the tail as an answer, we ask GPT4 to generate 15 possible candidate instance events with their event types for prompting.
We recruit another 10 annotators. The annotation is completed on our annotating system. Each annotator should rewrite correct data alongside at least the following standards:
1)
Rewrite and making them correct as high-probability knowledge, and do not consider low-probability situations.
2)
Rewrite and leading to the coherence of the whole graph, and there’s no semantic drift.
3)
Rewrite and making them consistent. Schema events and instance events require a clear distinction in hyper-hypo relation.
4)
Rewrite the target event making it only can be answered when considering the whole graph.
5)
Rewrite the instance events making them should be expressed independently without connective expressions such as "After ". This could incur information leakage.
We describe the detailed annotation process in the Appendix.
3.4 Question Adaptation
The last step is to construct questions of CEC and CRR in both schema and instance levels based on the annotated graphs. We use the schema part of annotation as the schema-level questions and the instance part as instance-level questions.
For schema-level CEC, we naturally use the queried event and other nodes except for the answer event as context to form a question. Then we use the answer event as the answer. We do similarly at the instance level. For CRR, we regard and as queried events and use the relation between them as the answer to form the schema-level question. For instance part, we adopt a similar way.
Finally, our CEC task is a 4-way multiple-choice task. The CRR is a 3-way multiple-choice task. In CRR, the choices for temporal, causal, and hierarchy relations are [Before, After, Vague], [Causes, IsResult, None], and [IsSubevent, HasSubevent, None] respectively. We show examples of both tasks in Figure 1. We report the number of each task and the average nodes and edges of in Figure 2.
@@ 3.5 Quality Inspection
After the annotation, to further guarantee the validation of , we recruit another group of human annotators to inspect the quality of . We sample 100 data for all tasks. We ask them to give two scores for each sample:
@@ 4.1 Evaluated LLMs
We evaluate 11 LLMs on event reasoning. For the open-source models, we evaluate their chat-version. We evaluate GPT4o, GPT4, GPT3.5. For the closed-source models, we utilize their official APIs to conduct performance evaluations. For the open-source models, we include Mistral-7B Jiang et al. (2023), Qwen2-7B Yang et al. (2024), Baichuan-2-7B Yang et al. (2023), Orca2-7B Mitra et al. (2023), Chatglm2-6B GLM et al. (2024), Internlm2-7B Cai et al. (2024), Llama2-7B Touvron et al. (2023), and Vicuna-7B Chiang et al. (2023). Without loss of generosity, we use the model names to refer to the chat versions in the rest of our paper. For all evaluated LLMs, we use the same prompt. We show all prompts in the Appendix.
@@ 5.1 How proficient abilities of event reasoning do LLMs have?
Model | S-CEC | I-CEC | S-CRR | I-CRR |
GPT4o | 68.93 | 66.60 | 62.05 | 62.04 |
GPT4 | 68.11 | 68.43 | 63.01 | 63.81 |
GPT3.5 | 65.43 | 64.77 | 54.52 | 43.95 |
Mistral-7B | 52.47 | 54.18 | 51.64 | 55.24 |
Qwen2-7B | 62.14 | 63.75 | 52.05 | 47.62 |
Baichuan2-7B | 52.88 | 29.94 | 51.64 | 45.31 |
Orca2-7B | 59.88 | 60.08 | 46.16 | 45.17 |
Chatglm2-6B | 55.76 | 30.96 | 52.47 | 49.66 |
Interlm2-7B | 65.84 | 62.12 | 48.63 | 57.28 |
Llama2-7B | 45.06 | 29.74 | 46.58 | 41.22 |
Vicuna-7b | 25.93 | 27.09 | 52.05 | 51.43 |
@@
We evaluate CEC and CRR at the instance level. We show the results of different relations in Figure 2 and detailed results in the Appendix. For CEC, GPT4 performs the best. Among all open-source LLMs, Qwen2 is the best while Internlm2 holds the second. Early models such as Baichuan2-7B, Chatglm2-6B, Llama2-7B, and Vicuna-7B fall in this task where they score under 40%. For CRR, GPT4 excels all other models as well. Among all open-source LLMs, Internlm2-7B and Mistral-7B performs in the first tie. There is no obvious difference in the performance of other models for CRR.
We show the average performance of instance-level CEC and CRR in columns I-CEC and I-CRR in Table 4. Overall, existing LLMs such as GPT4, and Qwen2-7B have CEC event reasoning abilities. However, even the strongest GPT4 can only achieve 68.43 (4-way multiple choice) and 63.81 (3-way multiple choice) accuracy in each task showing there’s much room for improvements of event reasoning.
@@
Comparing CEC to CRR, as relation-wise results shown in Figure 2 and average performances in columns I-CEC and I-CRR in Table 4, LLMs perform better for CEC than CRR (note that CEC is a 4-way multiple choice task while CRR is of 3-way). To compare, we compute the average scores of I-CEC and I-CRR on models achieving above 40% and 30%444Models under these scores may lack statistic significance.. We find I-CEC is much higher than I-CRR, with average scores 62.84 and 53.58. The results significantly suggest that CRR is harder and insifeciently solved than CEC. Existing pretraining and SFT datasets may be biased in paradigms.
We then analyze performances on different relations. We compute the average scores of relations on models achieving above 40% and 30% on average I-CEC and I-CRR. The I-CEC average scores of Temporal, Causal, Hierarchical are 50.11, 68.71, and 61.22 while in I-CRR the scores are 43.31, 58.36, and 52.17. With these results alongside scores shown in Figure 2, LLMs perform best in Causal relation. Then, Temporal relation is the worst. It indicates current training can enable LLMs to reason causality. It also trains the event hierarchy comprehension. However, temporal reasoning is the hardest. More methods should be established to handle this problem. That further indicates the imbalance training of different relations. Methods and datasets of balanced abilities on relations are needed. Transferring abilities of different relations could also be feasible Tao et al. (2023b).
This is a crucial finding. Chan et al. (2023) conduct causal event classification such as ECARE Du et al. (2022a), and relation reasoning such as MATRES Ning et al. (2018). They directly compare these two groups of results and conclude the gaps are merely from differences in relations. However, they ignore the difference in reasoning paradigms. Leveraging , with disentangling relations and formulations, we investigate event reasoning with less bias.
@@
We investigate the improvement trends of CEC and CRR. In Figure 3. When models have poor event reasoning abilities, their CEC performances lie around the balanced line showing no significant differences in tasks. With the development, the CEC improves much faster than CRR for all models. This investigation further appeals to the need for training in comprehensive and balanced event reasoning abilities.
CEC | CRR | |||
ET | Rel | ET | Rel | |
GPT4o | 65.53 | 40.23 | 69.84 | 51.24 |
GPT4 | 72.78 | 40.57 | 73.19 | 52.79 |
GPT3.5 | 10.05 | 15.58 | 16.19 | 28.63 |
Mistral-7B | 22.93 | 16.28 | 25.00 | 23.87 |
Qwen2-7B | 11.40 | 15.97 | 13.07 | 16.58 |
Model | Temporal | Causal | Hierarchical | w.t.S | w.o.S | |
GPT4o | 58.06 | 73.45 | 85.71 | 71.49 | 66.60 | 4.89 |
GPT4 | 61.29 | 75.86 | 77.92 | 72.51 | 68.43 | 4.08 |
GPT3.5 | 52.42 | 77.59 | 77.92 | 71.28 | 64.77 | 6.51 |
Mistral-7B | 45.16 | 70.34 | 74.03 | 64.56 | 54.18 | 10.38 |
Qwen2-7B | 61.29 | 78.97 | 77.92 | 74.34 | 63.75 | 10.59 |
Baichuan2-7B | 41.13 | 56.90 | 57.14 | 52.95 | 29.94 | 23.01 |
Orca2-7B | 53.23 | 77.59 | 76.62 | 71.28 | 60.08 | 11.20 |
Chatglm2-6B | 38.71 | 65.86 | 54.55 | 57.23 | 30.96 | 26.27 |
Interlm2-7B | 57.26 | 80.69 | 81.82 | 74.95 | 62.12 | 12.83 |
Llama2-7B | 36.29 | 45.86 | 49.35 | 43.99 | 29.74 | 14.25 |
Vicuna-7b | 27.42 | 28.62 | 27.27 | 28.11 | 27.09 | 1.02 |
Model | Temporal | Causal | Hierarchical | w.t.S | w.o.S | |
GPT4o | 56.45 | 73.04 | 75.65 | 69.25 | 62.04 | 7.21 |
GPT4 | 53.23 | 70.74 | 68.70 | 65.99 | 63.81 | 2.18 |
GPT3.5 | 44.09 | 59.22 | 47.83 | 53.61 | 43.95 | 9.66 |
Mistral-7B | 48.92 | 63.36 | 50.43 | 57.69 | 55.24 | 2.45 |
Qwen2-7B | 47.31 | 66.13 | 55.65 | 59.73 | 47.62 | 12.11 |
Baichuan2-7B | 46.24 | 58.29 | 40.87 | 52.52 | 45.31 | 7.21 |
Orca2-7B | 44.09 | 65.9 | 51.30 | 58.10 | 45.17 | 12.93 |
Chatglm2-6B | 47.31 | 53.00 | 50.43 | 51.16 | 49.66 | 1.50 |
Interlm2-7B | 53.23 | 70.28 | 72.17 | 66.26 | 57.28 | 8.98 |
Llama2-7B | 42.47 | 51.61 | 53.91 | 49.66 | 41.22 | 8.44 |
Vicuna-7b | 45.16 | 54.15 | 52.17 | 51.56 | 51.43 | 0.13 |